The sum of three numbers is 109. The first number is 7 less than the second and third number is four times the first. what are the numbers?
Answer by rfer(16322)

x+x-7+4(x-7)=109
x+x-7+4x-28=109
6x=109+7+28
6x=144
x=144/6
x=24 2nd
x-7=17 1st
4(x-7)=68 3rd